Q.5.The odometer of a car reads 57321.0 km when the clock shows the time 08:30 AM. What is the distance moved by the car, if at 08:50 AM, the odometer reading has changed to 57336.0 km? Calculate the speed of the car in km/min during this time. Express the speed in km/h also.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Distance = 57336.0 km – 57321 km =15 km Speed in km/min =15km/20 min = 3/4 km/min Speed in km/hr =15 km/1/3 hr =(15 x 3) km/hr =45 km/hr.